sinking fund payment. A company establishes a sinking fund to provide for the payment of a $100,000 debt maturing in 4 years. Find the amount of each annual deposit if interests 3% annum.
Answer #1

This question requires application of FV of annuity formula, according to which

[(1+r) -1 FV of Annuity = P P= Periodic Payment r=rate per period n = number of periods

(1 +0.03) - 1 100,000 = P*

100,000 = P * 4.183627

P = $23,902.70